# Leveraging Self-Supervised Audio-Visual Pretrained Models to Improve Vocoded Speech Intelligibility in Cochlear Implant Simulation

### v2

@@ I Introduction
Previous studies have confirmed that under quiet conditions, CI can effectively enhance the hearing capability of recipients, especially for speech recognition [97, 96, 29, 28]. However, it is also been reported that speech recognition performance degraded considerably when the target speech signals are distorted [6, 30, 76, 71]. In real-world scenarios, there are several distortion sources, including background noise, reverberation and interfering speech. To address speech distortion issues, a speech enhancement (SE) unit is usually adopted as a front-end processing unit in CI devices [31, 87]. Various techniques, including single-channel SE algorithms like spectral subtraction [85, 46], subspace methods [54], optimized gain functions [59], and commercial solutions [12, 13], have all been applied to improve CI performance. Furthermore, multi-microphone and beamforming approaches have been explored for SE in CI users, taking advantage of spatial filtering to better isolate speech signals from background noise [32, 44, 45, 81].
@@
In recent years, SE techniques have notably improved thanks to advances in machine learning algorithms. Notable examples include non-negative matrix factorization [91, 65], sparse coding [74, 79], compressive sensing [88], and robust principle component analysis [36]. More recently, further enhanced performance has been achieved using the strong regression capabilities of deep learning-based neural networks [55, 92, 86, 93, 98, 17, 34, 89, 94, 68, 69]. For these approaches, deep neural networks are often used as a mapping function to carry out enhancement filtering on noisy input to attain high-quality speech signals. Several extensions have been made to these deep-learning models. One direction is to use a more suitable objective function to train the SE system. In [47, 23, 48, 27, 22, 24, 34], speech metric-oriented objective functions are derived, which can be divided into two categories. The first category directly considers a particular metric to form the objective function, such as [47, 23, 48]. The second uses another neural network model to form the objective function, such as [27, 22, 24, 34]. Experimental results confirm that when a speech-metric oriented objective function is used, the SE system can be guided to achieve desirable output with optimal speech metric scores. In addition to designing more suitable objective functions, some researchers have attempted to incorporate information from other modalities as auxiliary inputs to the SE model, enabling exploitation of additional contextual information. Visual clues are one important modality that carries complementary information to speech signals during everyday communication. For example, the McGurk effect [61] demonstrates the cross-effect between the visualized mouth or lip shape and human hearing perception. From this viewpoint, numerous audio-visual multi-modal SE approaches, termed AVSE, have been proposed [25, 16, 33, 62, 41, 73, 63, 9, 8]. These studies clearly show that visual cues can successfully enhance the performance of audio-only speech enhancement (AOSE).
Developing an efficient AVSE system with limited training data is a crucial undertaking in real-world scenarios. Following [8], we address this challenge by exploiting self-supervised learning (SSL) in AVSE. The motivation for our work is the ability of SSL models to automatically label features and generate them based on knowledge of the input data. Specifically, instead of providing simple labels for each object as is the case for most supervised methods, SSL allows the extraction of more representative features such as the relative position of certain objects and their contextual relationship. This boosts performance by improving inference and reduces the amount of raw data required for pre-training [49, 14]. For example, the well-known Bidirectional Encoder Representations from Transformers (BERT) model produces contextual representations of languages generated using only a plain text corpus. BERT is a versatile AI pre-training model originally developed for Natural Language Processing (NLP). Compared to previous models using supervised methods, BERT has been shown to deliver substantial improvements in performance for a wide variety of tasks such as NLP and word recognition [43, 78, 37]. HuBERT, a BERT-derived model using hidden-units [35], has proven to be effective for denoising from an noise speech [38, 39]. In this paper, we develop a novel SSL-AVSE framework by leveraging AV-HuBERT [77], which is a HuBERT-based audio-visual model.
@@
In the past decade, deep learning-based SE techniques have been applied to CI systems [4, 50, 51, 42, 58], enabling more adaptive and robust processing methods tailored to the diverse auditory environments experienced by CI users. While the benefits of incorporating visual cues into the SE process are well-established, their specific advantages for CI devices have, to our knowledge, not yet been explored. In this study, we aim to evaluate the effectiveness of the proposed SSL-AVSE system for CI, with the implementation and user test scenario illustrated in Fig. 1. As shown in the figure, the SSL-AVSE system processes the combined noisy speech and video as inputs and outputs enhanced speech, which is then provided to a CI device. In this study, we employ a 16-channel speech vocoder to process enhanced utterances, simulating CI audio and calculating relevant performance metrics. Additionally, a listening test with 80 subjects is conducted to assess subjective speech quality and word intelligibility. Results indicate that our SSL-AVSE model achieves a maximum improvement of over 50% and 45% in speech quality and intelligibility compared to the original noisy signals, respectively, and a maximum improvement of over 20% and 40% compared to the AOSE system, respectively.
@@ II Methods
II-A Problem Formulation
Suppose we have a speech signal $s(t)$ and a noise signal $n(t)$. The noisy signal $x(t)$ can then be denoted as:
@@ II-A Problem Formulation
Next, the noisy signal $x(t)$ is transformed into the spectral feature $X$. The model predicts a ratio mask $M$ to extract the clean speech signal for the corresponding target speaker from $X$. $M$ is trained from our SSL-AVSE model, which uses the noisy signal $x(t)$ and additional visual cues $v(t)$ which is comprised of the lip image sequence. The enhanced speech can then be obtained by the following formula:
@@
where $\otimes$” indicates an element-wise multiplication. From the above, we can then use the enhanced representation, $\hat{S}$, to reconstruct the waveform signal $\hat{s}(t)$, also known as enhanced audio.
II-B Audio-Visual Speech Enhancement Networks
In this paper, we propose a novel SSL-AVSE framework, as shown in Fig. 2. Specifically, we combine the SE model with a Transformer-based AV-HuBERT network using self-attention mechanisms to provide enhanced performance.
II-B1 Data Preprocessing
The model takes the visual stream of detected lip images from the target speaker $v(t)$, the noisy speech $x(t)$ as input and outputs the enhanced speech $\hat{s}(t)$ for the target speaker while suppressing noise sources.
Audio Preprocessing. In this study, the noisy speech signal $x(t)$ is first transformed into spectrograms using the Short-Time Fourier Transform (STFT). Subsequently, the $log1p$ function ($log1p(z)$ = log(1+$z$)) is applied to these spectrograms to extract $log1p$ spectral features. It has been demonstrated in our prior research that these $log1p$ spectral features outperform conventional log power spectral features in terms of SE performance [39, 21]. As shown in Fig. 2, the noisy $log1p$ spectral features are multiplied by the estimated mask to produce enhanced speech, as shown in Eq. (2). The mask is generated by our SSL-AVSE system, which is based on AV-Hubert and takes the video signal, $v(t)$, along with the noisy speech signal, $x(t)$, as input.
Video Preprocessing. The video signal, $v(t)$, is comprised of the sequential images sampled at 25 fps and are cropped around the target speaker’s mouth region of interest (ROI). This is determined with a facial landmark detector consisting of a two-dimensional Face Alignment Network (FAN)[5]. Assuming a 3-second video clip at 25 fps, we can obtain a total of 75 lip images from the target speaker.
II-B2 Model
The representations of each transformer-encoder layer are denoted as $H^{l}$, where $10\leq l\leq L-$ 1 and $L$ is the number of layers. A trainable function $w(\cdot)$ is then applied to all of the layer representations as follows:
$H_{WS}=\sum_{l=0}^{L-1}w^{l}H^{l},$ (3)
@@ II-B2 Model
The features are then fed into the SE model consisting of a two-layer bidirectional long short-term memory (BLSTM) module sandwiched between two linear layers which performs forward and back propagation simultaneously. The output of the SE module is a soft mask and is multiplied by the magnitude of the noisy speech spectra. This is then compared with the clean speech spectra to determine the L1 (absolute) loss.
@@
While the video segment length used for learning SSL-AVSE is fixed, at the inference stage, our audiovisual extraction model can be applied to process videos of arbitrary length. This is done by applying a sliding window technique, which shifts the proposed window along the video segment until its entire length is covered.
@@
Figure 2: The proposed SSL-AVSE model includes AV-HuBERT and SE modules. The AV-HuBERT model consisting of multiple transformer layers is used to extract key features from noisy speech and lip images. An SE model then performs enhancement using them.
II-C CI Vocoded Speech
@@
We passed voice signals through vocoder to simulate CI sounds. These simulations were then played to normal hearing (NH) people to conduct an audiometric test [18, 75]. Compared with ordinary speech, vocoded speech is more difficult to understand by NH listeners due to the loss of spectral detail. Several studies have examined CI vocoder simulated speech on NH subjects in order to understand the associations between specific factors and CI users[80, 20, 19, 50]. Because accurate CI sounds are not always readily available, vocoder simulations can avoid the manifestation of patient-specific confounding factors, such as neural survival patterns [53]. Therefore, the CI vocoder can serve as an invaluable tool in related research.
@@
To simulate CI audio in this study, the tone vocoder was performed for an audio signal by using the procedure depicted in Fig. 3. As shown in the figure, there are four steps: (1) 16 Butterworth band-pass filters were used to process an input temporal sequence to produce bandpass signals. (2) For each band waveform, a full-wave rectification function was leveraged to smooth the signal and to generate the corresponding envelope wave. (3) We added a tonal signal to the envelope to produce the modulated band voice. (4) We summed all modulated voice and performed normalization operation to generate the vocoded speech, which has the identical root-mean-square value to the original input signal.
@@
Figure 3: Block diagram of our tone vocoder. The system consists of three modules: (1) A band-pass filter, (2) An envelope extractor and (3) a compressor. The results of each band are then summed and combined to form vocoded speech.
II-D Measures of Speech Intelligibility
Speech intelligibility is assessed through objective and subjective measures in scientific research. The three most widely used measures include the Perceptual Evaluation of Speech Quality (PESQ) [70] and the Short Time Objective Intelligibility (STOI) [82] for non-vocoded speech and the Normalized Covariance Metric (NCM) [57] for vocoded speech. PESQ is an algorithm designed to predict subjective opinion scores of a degraded audio sample. It provides numerical values ranging from -0.5 to 4.5. This measure is considered a full reference algorithm as it compares both the original and degraded signal [84]. STOI is another important measure of speech quality that is often used in SE settings. It is highly correlated with the intelligibility of degraded speech signals and takes into consideration the effect of non-linear processing on noisy speech. NCM is derived from multivariate normal distribution and is used to determine the correlation between multiple variables which are usually described as a matrix. Its dimension is dependent on the number of channels involved; since our vocoder simulator consists of 16 distinctive channels, a 16-dimensional NCM can be used to calculate their correlation coefficients. This value depends on the relationship between enhanced and clean speech; the greater the similarity the larger the NCM value. Orthogonal matrices, which consist of completely independent variables, have a correlation coefficient of 0, while congruent matrices, which consist of completely dependent variables, have a correlation coefficient of 1. To compare performance, we used PESQ and STOI scores to evaluate the SE results for normal speech. For CI-vocoded speech, we used NCM for performance comparison, as suggested in [7].
@@ II-D Measures of Speech Intelligibility
Subjective measures used in this study include overall speech quality and word intelligibility. The former is determined when subjects are asked to rate the quality of an entire utterance from 1 to 5. The higher the score, the better the perceived quality of the recorded sentence. The latter is based on whether subjects can clearly understand individual words in the given sentence. Since each utterance contains a total of 10 Chinese characters, a score of 1.0 (=10/10) indicates a complete understanding of each character in the given target sentence.
II-E Implementation Details
II-E1 Ablation Studies
To confirm the effectiveness of incorporating the pre-trained AV model, we built an AOSE system employing a BLSTM (Bidirectional Long-short Term Memory) model. In addition, we added ResNET-18 to process visual cues and combined it with the original one to form an AVSE model. Finally, we replace ResNET-18 with AV-HuBERT which allows for self-supervised learning.
II-E2 Model Training

### v3

@@ Abstract
Objective: Individuals with hearing impairments face challenges in their ability to comprehend speech, particularly in noisy environments. This study explores the effectiveness of audio-visual speech enhancement (AVSE) in improving the intelligibility of vocoded speech in cochlear implant (CI) simulations. Methods: We propose a speech enhancement framework called Self-Supervised Learning-based AVSE (SSL-AVSE), which uses visual cues such as lip and mouth movements along with corresponding speech. Features are extracted using the AV-HuBERT model and refined through a bidirectional LSTM. Experiments were conducted using the Taiwan Mandarin speech with video (TMSV) dataset. Results: Objective evaluations showed improvements in PESQ from 1.43 to 1.67 and in STOI from 0.70 to 0.74. NCM scores increased by up to 87.2% over the noisy baseline. Subjective listening tests further demonstrated maximum gains of 45.2% in speech quality and 51.9% in word intelligibility. Conclusion: SSL-AVSE consistently outperforms AOSE and conventional AVSE baselines. Listening tests with statistically significant results confirm its effectiveness. In addition to its strong performance, SSL-AVSE demonstrates cross-lingual generalization: although it was pretrained on English data, it performs effectively on Mandarin speech. This finding highlights the robustness of the features extracted by a pretrained foundation model and their applicability across languages. Significance: To the best of our knowledge, no prior work has explored the application of AVSE to CI simulations. This study provides the first evidence that incorporating visual information can significantly improve the intelligibility of vocoded speech in CI scenarios.
{IEEEkeywords}
Audio-visual speech enhancement, cochlear implants, self-supervised learning, cross-lingual generalization.
1 Introduction
@@
Previous studies have confirmed that under quiet conditions, CI can effectively enhance the hearing capability of recipients, especially for speech recognition [5, 11, 12, 13]. However, it has been reported that speech recognition performance degraded considerably when the target speech signals are distorted [14, 15, 16, 17]. In real-world scenarios, there are several distortion sources, including background noise, reverberation, and interfering speech. To address speech distortion issues, a speech enhancement (SE) unit is usually adopted as a front-end processing unit in CI devices [18, 19]. Various techniques, including single-channel SE algorithms like spectral subtraction [20, 21], subspace methods [22], optimized gain functions [23], and commercial solutions [24, 25], have all been applied to improve CI performance. Furthermore, multi-microphone and beamforming approaches have been explored for SE in CI users, taking advantage of spatial filtering to better isolate speech signals from background noise [26, 27, 28, 29].
@@
In recent years, SE techniques have improved significantly thanks to advances in machine learning algorithms. Notable examples include non-negative matrix factorization [30, 31], sparse coding [32, 33], compressive sensing [34], and robust principal component analysis [35]. More recently, further advancements have been achieved through the powerful regression capabilities of deep learning–based models [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. For these approaches, deep neural networks are often used as a mapping function to carry out enhancement filtering on noisy input to attain high-quality speech signals. Several extensions have been made to these deep-learning models. One direction is to use a more suitable objective function to train the SE system. In [53, 54, 55, 56, 57, 58, 42], speech metric-oriented objective functions are derived, which can be divided into two categories. The first category directly considers a particular metric to form the objective function, such as [53, 54, 55, 59]. The second uses another neural network model to form the objective function, such as [56, 57, 58, 42]. Experimental results confirm that when a speech-metric oriented objective function is used, the SE system can be guided to achieve desirable output with optimal speech metric scores. In addition to designing more suitable objective functions, some researchers have attempted to incorporate information from other modalities as auxiliary inputs to the SE model, enabling exploitation of additional contextual information. Visual clues are one important modality that carries complementary information to speech signals during everyday communication. Numerous audio-visual multi-modal SE approaches, termed AVSE, have been proposed [60, 61, 62, 63, 64, 65, 66, 67, 68, 69]. These studies clearly show that visual cues can successfully enhance the performance of audio-only speech enhancement (AOSE).
Developing an efficient AVSE system with limited training data is a critical challenge in real-world applications. Following [68], we address this issue by leveraging self-supervised learning (SSL) in AVSE. SSL models are trained by reconstructing the original input at the output, thereby learning to effectively analyze and resynthesize the data without requiring labels. Across numerous tasks, SSL has demonstrated its ability to extract more representative features, thereby boosting performance in downstream classification and regression tasks [70, 71]. For example, the well-known Bidirectional Encoder Representations from Transformers (BERT) model generates contextual language representations from large text corpora. As a versatile AI pretraining model originally developed for Natural Language Processing (NLP), BERT has demonstrated substantial performance improvements over previous supervised approaches across a wide range of tasks, including language understanding and speech recognition [72, 73, 74]. In speech processing, HuBERT—a BERT-derived model based on hidden units [75]—has shown strong effectiveness for the SE task [76, 77]. Building on these advances, we propose a novel SSL-AVSE framework that leverages AV-HuBERT [78], an audio-visual extension of HuBERT.
@@
In the past decade, deep learning-based SE techniques have been applied to CI systems [79, 80, 81, 82, 83, 84], enabling more adaptive and robust processing methods tailored to the diverse auditory environments experienced by CI users. While the benefits of incorporating visual cues into the SE process are well-established, their specific advantages for CI devices have, to our knowledge, not yet been explored. In this study, we aim to evaluate the effectiveness of the proposed SSL-AVSE system for CI, with the implementation and user test scenario illustrated in Fig. 1. As shown in the figure, the SSL-AVSE system processes the combined noisy speech and video as inputs and outputs enhanced speech, which is then provided to a CI device. In this study, we employ a 16-channel speech vocoder to process enhanced utterances, simulating CI audio and calculating relevant performance metrics. Additionally, a listening test with 80 subjects is conducted to assess subjective speech quality and word intelligibility. Results indicate that the proposed SSL-AVSE model achieves maximum improvements of over 45% in speech quality and 50% in intelligibility compared with the original noisy signals, and over 25% and 45% improvements, respectively, compared with the baseline AVSE system.
@@ 2 Methods
2.1 Problem Formulation
Given a speech signal $s(t)$ and a noise signal $n(t)$, the noisy signal $x(t)$ can then be denoted as:
@@ 2.1 Problem Formulation
The noisy signal $x(t)$ is transformed into the spectral feature $X$. The model predicts a ratio mask $M$ to extract the clean speech signal for the corresponding target speaker from $X$. $M$ is estimated from our SSL-AVSE model, which uses the noisy signal $x(t)$ and additional visual cues $v(t)$ composed of the lip image sequence. The enhanced speech can then be obtained by the following formula:
@@
where $\otimes$” indicates an element-wise multiplication. From the above, we can then use the enhanced representation, $\hat{S}$, to reconstruct the waveform signal $\hat{s}(t)$, also known as enhanced speech.
2.2 Audio-Visual Speech Enhancement Networks
In this paper, we propose a novel SSL-AVSE framework, as illustrated in Fig. 2. Specifically, SSL-AVSE integrates a Transformer-based AV-HuBERT network, a pretrained audio-visual foundation model, with the SE model.
2.2.1 Data Preprocessing
The model takes the visual stream of detected lip images from the target speaker $v(t)$, the noisy speech $x(t)$ as input, and outputs the enhanced speech $\hat{s}(t)$ for the target speaker while suppressing noise signals.
Audio Preprocessing. In this study, the noisy speech signal $x(t)$ is first transformed into spectrograms using the Short-Time Fourier Transform (STFT) with an FFT size of 512, a window length of 400, and a hop size of 160. Subsequently, the $log1p$ function ($log1p(z)$ = log(1+$z$)) is applied to these spectrograms to extract $log1p$ spectral features. It has been demonstrated in our prior research that these $log1p$ spectral features outperform conventional log power spectral features in terms of SE performance [77, 85]. As shown in Fig. 2, the noisy $log1p$ spectral features are multiplied by the estimated mask to produce enhanced speech. The mask is generated by the SSL-AVSE system, which is based on AV-Hubert and takes the video signal, $v(t)$, along with the noisy speech signal, $x(t)$, as input.
Video Preprocessing. The video signal, $v(t)$, is comprised of sequential images sampled at 50 frames per second (fps), cropped around the target speaker’s mouth region of interest (ROI). The ROI is detected using a facial landmark detector based on a two-dimensional Face Alignment Network (FAN) [86], center-cropped to 88 × 88 pixels, and normalized using [0, 255] scaling followed by standardization with mean 0 and standard deviation 1.
2.2.2 Model
The representations of each Transformer-encoder layer are denoted as $H^{l}$, where $0\leq l\leq L-1$ and $L$ is the number of layers. A trainable function $w(\cdot)$ is then applied to all of the layer representations as follows:
$H_{WS}=\sum_{l=0}^{L-1}w^{l}H^{l},\$(3)
@@ 2.2.2 Model
The extracted features are then passed to the SE model, which consists of a two-layer bidirectional long short-term memory (BLSTM) module positioned between two linear layers. The output of the SE module is a soft mask and is multiplied by the magnitude of the noisy speech spectra. This is then compared with the clean speech spectra to determine the L1 (absolute) loss.
@@
While the video segment length used for learning SSL-AVSE is fixed, at the inference stage, our audio-visual extraction model can be applied to process videos of arbitrary length. This is done by applying a sliding window technique, which shifts the proposed window along the video segment until its entire length is covered.
@@
Figure 2: The proposed SSL-AVSE model includes AV-HuBERT and SE modules. The AV-HuBERT model, consisting of multiple Transformer layers, is used to extract key features from noisy speech and lip images. An SE model then performs enhancement using them.
2.3 CI Vocoded Speech
@@
We passed voice signals through a vocoder to simulate CI sounds. These simulations were then played to normal hearing (NH) people to conduct a listening test [87, 88]. Compared with ordinary speech, vocoded speech is more difficult to understand by NH listeners due to the loss of spectral detail. Several studies have examined CI vocoder simulated speech on NH subjects in order to understand the associations between specific factors and CI users [89, 90, 91, 80]. Because accurate CI sounds are not always readily available, vocoder simulations can avoid the manifestation of patient-specific confounding factors, such as neural survival patterns [92]. Therefore, the CI vocoder can serve as an invaluable tool in related research.
@@
To simulate CI audio in this study, a tone vocoder was used to process the speech signals following the procedure illustrated in Fig. 3. As shown in the figure, there are four steps: (1) 16 Butterworth band-pass filters were used to process an input temporal sequence to produce bandpass signals. (2) For each band waveform, a full-wave rectification function was leveraged to smooth the signal and to generate the corresponding envelope wave. (3) We added a tonal signal to the envelope to produce the modulated band voice. (4) We summed all modulated voice and performed a normalization operation to generate the vocoded speech, which has the identical root-mean-square value to the original input signal.
@@
Figure 3: Block diagram of the tone vocoder used in this study. The system consists of a set of band-pass filters, an envelope extractor, and sinewave carriers. The results of each band are then summed and combined to form vocoded speech.
2.4 Measures of Speech Quality and Intelligibility
In this study, we employed four objective metrics to evaluate both non-vocoded and vocoded speech, namely: (1) for non-vocoded speech: the Perceptual Evaluation of Speech Quality (PESQ) [93], the Short Time Objective Intelligibility (STOI) [94], and the Levenshtein Phone Similarity (LPS) [95]; for vocoded speech: the Normalized Covariance Metric (NCM) [96]. PESQ is a metric designed to predict subjective opinion scores of degraded speech samples, providing numerical values ranging from –0.5 to 4.5. To compute PESQ, the distorted speech sample must be paired with its corresponding clean reference signal, making PESQ an intrusive approach [97]. STOI is a widely used measure of speech intelligibility in SE tasks. It has been shown to correlate well with the intelligibility of degraded speech signals and accounts for the effects of non-linear processing on noisy speech [94]. STOI scores range from 0 to 1. LPS measures the Levenshtein distance between phoneme sequences extracted from the enhanced and clean speech. Higher values denote stronger preservation of the intended phone sequence and are useful for detecting hallucinated [98]. NCM is a speech transmission index (TI)-related metric, estimated from the covariance of the envelopes between clean and processed signals. Prior studies have shown that NCM correlates well with the intelligibility of vocoded speech, largely because its calculation resembles CI processing strategies—both rely on envelope information across multiple frequency bands while discarding fine-structure details. NCM scores range from 0 to 1, with higher values indicating better intelligibility. Further details on the NCM measure can be found in [96]. To evaluate the effectiveness of the proposed SSL-AVSE, we used PESQ, STOI, and LPS to assess the SE results for non-vocoded speech. For CI-vocoded speech, performance was measured using NCM, as recommended in [99].
@@ 2.4 Measures of Speech Quality and Intelligibility
Subjective measures used in this study include overall speech quality and word intelligibility. The former was evaluated following the ITU-T P.808 protocol [100], where subjects were asked to rate the quality of an entire utterance on a 5-point scale. The higher the score, the better the perceived quality of the recorded sentence. The latter is based on whether subjects could clearly understand individual words in the given sentence. Since each utterance contains a total of 10 Chinese characters, a score of 1.0 (= 10/10) indicates a complete understanding of each character in the given target sentence.
2.5 Implementation Details
2.5.1 Ablation Studies
To validate the effectiveness of incorporating a pretrained AV foundation model, we first built an AOSE system using a BLSTM network. We then added a ResNet-18 module to process visual cues and integrated it with the AOSE system to establish a baseline AVSE model.
2.5.2 Model Training